OpenStudy (anonymous):

simplify the expression...... 15+ (-7)

OpenStudy (anonymous):

Is that a negative 7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got -2

OpenStudy (anonymous):

can u explain how u did it

Mehek (mehek14):

how do you get that???

Mehek (mehek14):

15 + (-7) = 15 - 7 = 8

OpenStudy (anonymous):

Oh so i did that wrong

Mehek (mehek14):

yes

OpenStudy (anonymous):

can someone help me!!!

Mehek (mehek14):

@miss.haleyyy I just did....

OpenStudy (anonymous):

Yeah he just told the answer

Mehek (mehek14):

she*

OpenStudy (anonymous):

sorry i meant she

OpenStudy (anonymous):

thxx both of u

OpenStudy (anonymous):

lol